Suppose that the salary range for recent college graduates with a bachelor's degree in economics is $30,000 to $50,000, with 25 percent of jobs offering $30,000 per year, 50 percent offering $40,000 per year and 25 percent offering $50,000 per year and that in all other respects, the jobs are equally satisfying. Assume that in this market, a job offer remains open for only a short time so that continuing to search requires an applicant to reject any current job offer. The expected starting salary for a college graduate with a bachelor's degree in economics is ______.

Select one:

a. $30,000

b. $40,000

c. $45,000

d. $50,000
Answer #1

The expected salary for a college graduate with a bachelor's degree in economics is calculated as weighted average of all payoffs, using probability of each payoff as its weight.

Thus, here the expected salary =

(25%) * 30,000 + (50%) * 40,000 + (25%) * 50,000

= 0.25 * 30,000 + 0.5 * 40,000+ 0.25 * 50,000

=7500 + 20,000 + 12,500 = 40,000.

Answer is (b) $40,000.